# Adaptive Interactive Navigation of Quadruped Robots using Large Language Models

### v1

@@ IV-D Ablation Studies
To further investigate the effectiveness of the proposed method, we conduct ablation study to compare AINav with the following four variants. 1) AINav-Single, which only generates a single plan with LLM in the task planning module; 2) AINav-Skill, which directly generates skills without incorporating object-centric reasoning; 3) AINav-No, which removes the replanning module; 4) AINav-F, which only replans when a failure is encountered. We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building, which need interaction for task completion, to compare the success rate and overall time. The results of ablation studies are presented in Fig. 6. AINav-Single exhibits low success rates, as the single plan generated by the LLM is often flawed or inefficient due to inherent uncertainties of LLM, leading to task failures. Benefiting from the tree structure that enhances the robustness and planning quality, AINav-Skill achieves a reasonable success rate. However, it lacks a comprehensive understanding of object utilization in complex scenarios, leading to task decomposition plans inexecutable. AINav-No performs poorly because it lacks a replanning phase, making it unable to adapt to new observations. Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, often resulting in excessive time spent and task timeouts. By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving higher success rates and improved time efficiency.
IV-E Performance of Primitive Skills
@@
We visualize the performance of various skills in Fig. 7. In Fig. 7(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks. Fig. 7(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment. In Fig. 7(c) and Fig. 7(d), we showcase the robot pushes boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy. The box-pushing process typically consists of three steps: first, the robot approaches the box and positions itself so that the box is between the robot and the target position, preparing for the subsequent pushing action; second, the robot pushes the box toward the target point; and finally, the robot fine-tunes the box’s position to complete the task. These versatile skills collectively form a comprehensive skill library that enables the robot to handle various sub-tasks in complex navigation scenarios.
V Experiment
Figure 8: Box-pushing policy in different configurations.
(a) Different scenarios for real-world experiments.
(b) Box obstruction
(c) Box utilization: indoor
(d) Box utilization: outdoor
Figure 9: Visualization of real-world experiments. (a) shows various scenarios in real-world experiments. (b), (c) and (d) illustrate case studies for each scenario, in which the top-left subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
We also conduct real-world experiments to validate the performance of AINav in practical situations. We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection. We evaluate the effectiveness of our method across three interactive navigation scenarios, as illustrated in Fig. 9(a). The first scenario involves an indoor obstacle obstruction setting, simulating a navigation task in a narrow corridor blocked by movable obstacles, where the robot needs to remove the box to reach the target. The second scenario, also set in a confined indoor corridor, featured a fixed $0.35$m-height platform obstructing the path. Since we assume the maximum climbing height for the robot is $0.3$ meters, the robot needs to reason how to traverse heights that exceed this capability limit by interacting with objects. The third scenario is box utilization in an outdoor environment, setting with fixed obstacles blocking the path and $0.45$m height, presenting a more difficult challenge with a different layout. Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain accurate object descriptions, thereby mitigating the risk of task failure due to perception errors.
Fig. 9(b) demonstrates the experiment on obstacle obstruction. Initially, the robot detects a blocking box and navigates around it using local point cloud information and its navigation policy to approach the box. Subsequently, a pushing skill is employed to move the obstacle aside, creating a path to the target. Fig. 9(c) illustrates how the robot uses a box to overcome a higher obstacle. After visually detecting the box, the task planning module decides to push the box next to the high platform as a step, enabling the robot to climb over the challenging obstacle and reach the target. Fig. 9(d) shows the robot completing the interactive navigation task in the outdoor scenario, employing task decomposition and skill execution to achieve the target, overcoming challenges that traditional navigation methods could not address. Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction. These above real-world experiments demonstrate our method’s capability to complete navigation tasks in challenging environments. Further details are available in our supplementary video. We conduct $5$ repeated experiments in each scenarios, and the success rates are $100\%$ ($/5$ 5), $80\%$ ($4/5$), and $80\%$ ($4/5$), respectively, highlighting the robustness and superiority of AINav in various interactive tasks and both indoor and outdoor settings.
To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig. 8. We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement. This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
VI Conclusion
We introduce an LLM-based adaptive interactive navigation approach using a quadruped robot, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals. Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning. The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained. Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal. Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations. This instructive framework has the potential to enhance robotic navigation capabilities, offering an effective and computationally efficient solution for interactive navigation in challenging environments. Future directions encompass developing more generalizable interaction skills to accommodate broader tasks and scenarios, and achieving tighter integration between LLM/VLM and versatile control policies for more intelligent embodied planning and control.

### v2

@@ V-D Ablation Studies
VI Real-World Experiment
We conduct real-world experiments to validate the performance of AINav in practical situations. We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection. We evaluate the effectiveness of AINav across five interactive navigation scenarios, as illustrated in Fig.˜10 (a). The first three scenarios are conducted in indoor environments: obstacle traversal, stair building, and corridor obstruction. In obstacle traversal, the robot is required to utilize a box as a stepping stone to surmount a 0.350.35m obstacle. The stair-building task requires the robot to arrange boxes of two different heights to construct a temporary staircase, enabling it to ascend a 0.500.50m platform. In corridor obstruction, two boxes obstruct a narrow passageway, one of which is heavy to push. The robot must proactively interact with these obstacles to create a navigable path and replan its strategy upon encountering the fixed box. The final two scenarios are set in outdoor environments, where the robot needs to utilize a box to traverse a 0.450.45m obstacle on the concrete floor or grass. We assume the quadruped’s maximum climbing height is 0.300.30m. We benchmark our method against the RoboTool baseline, which performs reasonably well in our simulations and has been validated through real-world experiments [xu2023creative]. We adopt the SLAM Toolbox package to obtain the robot’s real-time pose. Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain object pose and dimension, thereby eliminating the risk of confounding factor detection errors.
Figure 9: Box-pushing policy in different configurations. Each row demonstrates the pushing strategy under different initial box positions: left-shifted (top), centered (middle), and right-shifted (bottom).
Figure 10: Visualization of real-world experiments. (a) shows various scenarios in real-world experiments, where green stars indicate the navigation goal. Highlighted boxes indicate movable objects, with green markers denoting light movable boxes and red markers denoting heavy boxes. (b)-(f) illustrate case studies for each scenario, in which the leftmost subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
TABLE III: Metrics comparison in real-world experiments
Scenario RoboTool AINav
SR PT (s) ET (s) TLS (m) SR PT (s) ET (s) TLS (m)
Obstacle traversal $20\%$ 16.4 40.2 6.8 80% 11.8 38.4 8.1
Stair building $0\%$ - - - 20% 16.5 71.6 9.3
Corridor obstruction $40\%$ 15.9 58.4 6.9 80% 18.1 53.7 6.4
Urban climbing $60\%$ 14.3 45.2 7.0 80% 8.4 37.6 6.6
Field climbing 60% 19.2 62.4 12.3 60% 11.3 60.4 9.4
Average success rate $36\%$ 64%
Fig.˜10 (b) showcases AINav’s capacity for dynamic replanning and creative tool utilization. Initially, the robot attempts a direct pushing strategy to withdraw a box blocking its path. Upon discovering that the box (red highlighted in Fig.˜10 (a)) is difficult to push, AINav replans and pushes a nearby smaller box as a support structure to scale the obstacle. This adaptation from unpredictable environmental information underscores AINav’s ability to traverse impassable barriers through tool use and robust strategic replanning. Fig.˜10 (c) showcases AINav’s advanced tool utilization capability. The robot constructs a makeshift staircase from boxes of varying heights to ascend a 0.5m platform, which is twice the height of the robot itself, showing AINav’s long-horizon reasoning and planning abilities. Although the task planner consistently generates viable solutions, the relatively low success rate is primarily attributed to execution failures, particularly during the challenging task of sequential climbing on non-fixed boxes. Addressing this execution limitation will be a key focus of our future work. Fig.˜10 (d) demonstrates the experiment on corridor obstruction. After an initial attempt to push one of two obstructive boxes fails, the robot strategically interacts with the second box, clearing a navigable path to its destination. Fig.˜10 (e) and (f) show the robot completing the interactive navigation task in the outdoor scenarios, employing task decomposition and skill execution to achieve the navigation goal, overcoming challenges that traditional navigation methods could not address. Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction. These real-world experiments demonstrate AINav’s capability to perform interactive navigation tasks across diverse scenarios, highlighting its proficiency in both complex tool utilization and dynamic replanning in response to new environmental information. Further details are available in our supplementary video. Table˜III presents the metric comparison across different scenarios in $5$ repeated experiments, where AINav achieves a higher success rate, less planning time, and execution time, highlighting the effectiveness and adaptivity of AINav in various tasks and both indoor and outdoor settings.
To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig.˜9. We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement. This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
VII Conclusion
We introduce an LLM-based adaptive interactive navigation approach, AINav, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals. Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning. The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained. Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal. Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations. AINav has the potential to enhance robotic navigation capabilities, offering an effective and efficient solution for interactive navigation in challenging environments.
Future work will work on tighter integration between foundation models and versatile motion capabilities for real-time embodied planning and control. We plan to leverage visual foundation models to identify task-related objects and extract corresponding properties, broadening the practical applicability of AINav to real-world scenarios. Moreover, we will expand the variety of primitive skills and enhance their robustness and interior seamless transition, ultimately aiming to extend our framework to a diverse range of robotic platforms and complex, long-horizon tasks.
-A Prompts
-A1 Object-Level Proposer
You are a quadruped robot on the ground in a 3D world. Your goal is to navigate to a specific point in the 3D space. Your navigation goal is [goal point], and your current scene understanding is [scene understanding]. There are several objects in the scene that you may utilize. We use two parameters, position and size, to represent the location and size of an object, respectively. Each object’s position is represented by a 3D vector [x, y, z]. Each object’s size is represented as a 3D vector [length, width, height]. [object description]. You have a skill library containing the following skills and corresponding parameters: [skill library]. You must follow these constraints: [constraints]. Give me five different abstract plans for using objects to help you complete navigation tasks. The plan must include which objects you need to use, the sequence you use the objects, how to use the objects. You must analyze the problem step by step and show the thinking process. You must follow the following answer template:
[begin of plan]
Plan1: I need to use [objects]. First, … Second …
Plan2: …
[end of plan]
-A2 Skill-Level Proposer
You are a quadruped robot on the ground. Your goal is to navigate to a specific point in the 3D space. Your navigation goal is [goal point]. You have a skill library containing the following skills and corresponding parameters: [skill library]. Here are some abstract plans: [object-level plans], you need to generate detailed plans according to each abstract plan. Each step in the detailed plan consists of a skill with an abstract position like ’walk-to(’abstract position’)’.
You must follow the following answer template:
[begin of plan]
Plan1: [
(’step1’,’<skill>’),
[end of plan]
-A3 Evaluator
You are a quadruped robot on the ground in a 3D world. Your goal is to navigate to a specific point in the 3D space. Your navigation goal is [goal point], and your current scene understanding is [scene understanding] There are several objects in the scene that you may utilize: [object description]. You have a skill library containing the following skills and corresponding parameters: [skill library]. Here are some detailed plans to accomplish the navigation task: … You need to evaluate the reward of each step in the plan based on its contribution to task completion, assigning a value between 0 and 1. Specifically, this involves simulating whether future steps can reach the goal under the current step to assess its impact. If the step does not satisfy the given constraints, the reward is 0. Constraints you must follow: [constraints]. At the end of your response, reply to me with the following answer template:
[begin of evaluation]
Plan1: [
(’step1’, ’<skill>’, ’reward: …’),
[end of evaluation]
-A4 Parameter Calculation
The current plan is…, you need to calculate the 3D coordinates [x, y, z] of the abstract position in the plan based on the robot pose and the updated object information. The robot pose is [robot pose]. The object information is [object description]. You need to consider the spatial relationship between objects to obtain a reasonable position. You must calculate the position along each dimension step by step.
-A5 Advisor
You are a quadruped robot on the ground in a 3D world. Your goal is to navigate to a specific point in the 3D space. Your navigation goal is [goal point]. There are several objects in the scene that you may utilize: [object description]. Your current plan is: … Now, you have a new observation interpretation of the environment: … You need to determine whether to replan to modify your current plan based on the current plan, this new environmental observation, and the following criteria. You must analyze whether to reply, provide me with the reason, and respond with "Yes" or "No".
Replanning criteria:
* If the new observation is an execution failure of the current plan, then replanning is necessary.
* If the new observation is a new object, you need to evaluate how this new object might help complete your task. If using it results in a more effective plan than your current one, you need to replan.
* If the new observation is a revaluation of a previously known object, you need to determine whether this new information impacts your current plan. For example, if your plan requires climbing onto a box, but new observations show that the box is too high to climb, you need to replan.
-A6 Arborist
You are a quadruped robot on the ground in a 3D world. Your goal is to navigate to a specific point in the 3D space. Your navigation goal is [goal point]. There are several objects in the scene that you may utilize: [object description]. Constraints you must follow: [constraints]. Your current plan is: … Now, you have a new observation of the environment: …. Here are replanning suggestions by the advisor: … You must make adjustments to your current plans based on the suggestions provided by the advisor, such as expanding new skills for new objects or pruning infeasible skills.
-B Reward Design
[…]
TABLE IV: Definition of symbols.
Symbol Description
$\mathbf{q}_{j}$ Joint positions
$\dot{\mathbf{q}_{j}}$ Joint velocities
$\ddot{\mathbf{q}_{j}}$ Joint accelerations
$\mathbf{q}^{*}_{j}$} Target joint positions
$\mathbf{q}^{*,pre}_{j}$} Last target joint positions
$\boldsymbol{\tau}_{j}$ Joint torques
$\mathbf{v}_{b}$ Base linear velocity
$\boldsymbol{\omega}_{b}$ Base angular velocity
$\boldsymbol{x}_{b}$ Base position
$\boldsymbol{\theta}_{b}$ Base heading
$\boldsymbol{x}_{o}$ Object pose
$\mathbf{v}^{*}_{b}$ Commanded base linear velocity
$\boldsymbol{\omega}^{*}_{b}$ Commanded base angular velocity
$\boldsymbol{x}^{*}_{b}$ Commanded base position
$\boldsymbol{x}^{*}_{o}$ Commanded object pose
$n_{c}$ Number of collisions
$\mathbf{t}_{air,f}$ Feet air time
TABLE V: Reward terms in walking skill.
Reward terms Definition Weight
Linear velocity tracking $\phi(\mathbf{v}^{*}_{b,xy}-\mathbf{v}_{b,xy})$ 11
Angular velocity tracking $\phi(\boldsymbol{\omega}^{*}_{b,z}-\boldsymbol{\omega}_{b,z})$ $0.5$
Linear velocity penalty $\mathbf{v}_{b,z}^{2}$ 22
Angular velocity penalty $\boldsymbol{\omega}_{b,xy}||^{2}$ $0.05$
Joint torques $\boldsymbol{\tau}_{j}||^{2}$ $0.00001$
Joint accelerations $\ddot{\mathbf{q}_{j}}||^{2}$ $2.5*10^{-7}$
Action rate $\mathbf{q}^{*}_{j}-\mathbf{q}^{*,pre}_{j}||^{2}$ $0.01$
Collisions $n_{collision}$ 11
Feet air time $\sum_{f=0}^{4}(\mathbf{t}_{air,f}-0.5)$ $0.125$
TABLE VI: Reward terms in climbing skill.
Reward terms Definition Weight
[…]
Move direction $\text{cos}\langle\mathbf{v}_{b},\boldsymbol{x}^{*}_{b}-\boldsymbol{x}_{b}\rangle$ 22
Linear velocity penalty $\mathbf{v}_{b,z}^{2}$ 22
Angular velocity penalty $\boldsymbol{\omega}_{b,xy}||^{2}$ $0.05$
Joint torques $\boldsymbol{\tau}_{j}||^{2}$ $0.00001$
Joint accelerations $\ddot{\mathbf{q}_{j}}||^{2}$ $2.5*10^{-7}$
Action rate $\mathbf{q}^{*}_{j}-\mathbf{q}^{*,pre}_{j}||^{2}$ $0.01$
Collisions $n_{collision}$ 11
Feet air time $\sum_{f=0}^{4}(\mathbf{t}_{air,f}-0.5)$ $0.125$
TABLE VII: Reward terms in navigation skill.
Reward terms Definition Weight
[…]
Negative x-velocity penalty $\max(\mathbf{v}_{b,x},0)$ 22
Collisions $n_{collision}$ $10$
TABLE VIII: Reward terms in pushing skill.
Reward terms Definition Weight
[…]
Object heading tracking $\boldsymbol{x}^{*}_{o,h}-\boldsymbol{x}_{o,h}|| 11$
Negative x-velocity penalty $\max(\mathbf{v}_{b,x},0)$ 11
Face to object $\text{cos}\langle\mathbf{\theta}_{b},\boldsymbol{x}_{o}-\boldsymbol{x}_{b}\rangle$ 22
-C Performance of Primitive Skills
@@ VII Conclusion
We visualize the performance of various skills in Fig.˜11. In Fig.˜11(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks. Fig.˜11(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment. In Fig.˜11(c) and Fig.˜11(d), we showcase the robot pushing boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy. These versatile skills collectively form a comprehensive skill library that enables the robot to handle various sub-tasks in complex navigation scenarios.
-D Case Visualization with Concrete Prompts
We provide a detailed process for the obstacle traversal scenario in real-world experiments, including prompts and outputs for all LLM-based modules. This covers the initial task decomposition and the replanning stage in case of failure. Due to space limitations, we only present the most informative content in prompt and LLM output.
Figure 12: Visualization of prompts and output of each module.